A new project will cost $100,000. It will yield new sales revenues of $496,000, but will increase annual variable costs by $416,000 and annual fixed costs by $15,000. The project will also require initial investment of $22,000 in net working capital. The project will last for three years, and depreciation will be straight-line to zero. Interest expense will be $5,000. Given that the required rate of return on this project is 18%, and the marginal corporate tax rate is 40%, what is the annual operating cash flow from this project? Ignore the half-year rule.
Answer #1

After-tax income =(496000-416000-15000)*(1-0.40) =39000

Depreciation tax shield =(100000/3)*40% =13333.33

OCF =52333.33